# Skewed score: a statistical framework to assess autograders

### v1

@@ Abstract
The evaluation of large language model (LLM) outputs is increasingly performed by other LLMs, a setup commonly known as “LLM-as-a-judge”, or autograders. While autograders offer a scalable alternative to human evaluation, they have shown mixed reliability and may exhibit systematic biases, depending on response type, scoring methodology, domain specificity, and other factors. In this paper we propose a statistical framework based on Bayesian generalised linear models (GLMs) that enables researchers to simultaneously assess their autograders while also addressing their primary research questions (e.g., LLM evaluation). Our approach models evaluation outcomes (e.g., scores or pairwise preferences) as a function of properties of the grader (e.g., human vs. autograder) and the evaluated item (e.g., response length or the LLM that generated it), allowing for explicit quantification of scoring differences and potential biases within a unified framework. In addition, our method can be used to augment traditional reliability metrics such as inter-rater agreement, by providing uncertainty estimates and clarifying the source of disagreement. Overall, this approach contributes to more robust and interpretable use of autograders in LLM evaluation, enabling both performance analysis and bias detection.
@@ Introduction
Large language models (LLMs) are being deployed across a wide range of tasks (e.g., code writing, question answering, decision support, autonomous tool use like calling APIs and retrieving documents). To understand their capabilities and ensure reliable performance and safety, LLM outputs need to be continuously and systematically evaluated. Given the high volume and increasing complexity of these evaluations, automated approaches have become common. Unsurprisingly, LLMs themselves have emerged as candidates for this task, a practice referred to as “LLM-as-a-judge” [Zheng et al., 2023; Liu et al., 2023; Chiang and Lee, 2023; Fu et al., 2024; Bavaresco et al., 2025; Wang et al., 2025] or as autograding [Schneider et al., 2023; Guan et al., 2024]. For brevity, we here refer to these models as autograders.
@@
Recent studies suggest that autograders may indeed exhibit systematic biases. One such bias is self-bias, where LLM-based graders tend to assign higher scores to responses generated by the same LLM [Panickssery et al., 2024; Liu et al., 2024b], or more broadly to machine-generated content over human-written responses [Liu et al., 2023]. Another common issue is length bias, where longer answers are preferred regardless of their actual quality [Zheng et al., 2023; Dubois et al., 2024]. Additional biases include preferences for certain writing styles, answer structures, or keyword presence [e.g., Koo et al., 2024; Wang et al., 2024; Stureborg et al., 2024; Wu and Aji, 2025]. Moreover, graders may exhibit intransitive preferences (e.g., preferring A over B, B over C, yet C over A), revealing inconsistencies that cannot be captured by standard evaluation metrics [Xu et al., 2025].
@@
Existing evaluation methods are limited in their ability to isolate and quantify such effects and inconsistencies. They do not explain why disagreement arises, do not account for how grader behaviour may vary with grader identity, model characteristics, or context, and do not quantify uncertainty in these effects. While recent work has started employing statistical modeling (e.g., logistic regression for length bias [Dubois et al., 2024]), these approaches typically remain narrowly tailored to specific biases and do not offer an integrated evaluation framework. In this paper, we introduce a Bayesian generalised linear modeling framework that complements existing metrics and provides an interpretable, extensible, and statistically grounded view of grader behaviour.
@@
In the following sections, we begin by presenting the general framework. We then walk through a series of common questions to demonstrate how the framework can be used to answer them. We begin by explaining how autograder scores can be compared to human scores (Question 1.1), and how this comparison can be integrated into an LLM evaluation analysis (Question 1.2). We then gradually increase the complexity of the setup and show how different questions about autograder quality can be addressed within this integrated framework. Specifically, we show how to assess whether the graders are biased towards certain models being evaluated (Question 2) and how to quantify individual differences in the case of multiple graders (Question 3). We also show how to analyse item-level patterns (Question 4), which allows us to compute inter-rater agreement metrics with quantification of uncertainty and, crucially, to identify whether disagreement stems from noise or systematic bias (something that traditional metrics cannot disentangle). Finally, we show how this framework can be applied to a setting in which LLM outputs are compared directly to one another (as opposed to being assigned absolute scores), how to quantify intransitive (e.g., cyclic) preferences and how to assess whether graders have biases towards longer formats (Question 5). To provide researchers with a practical reference for applying the framework, we summarise the evaluation questions alongside how each can be formalised within our approach in table 1.
@@
We propose using generalised linear models (GLMs) as a practical and flexible approach for analysing autograder performance during LLM evaluations. GLMs extend standard linear regression by allowing the outcome variable to follow a wider range of distributions, making them suitable for many types of data beyond normally distributed outcomes. At their core, GLMs express the outcome as a function of a linear predictor, which is then transformed by a link function appropriate to the distribution of the outcome. The linear predictor is a mathematical expression that combines input variables in a weighted sum:
Linear $\text{Linear predictor}=\beta_{0}+\beta_{1}X_{1}+\beta_{2}X_{2}+...+\beta_{n}X_{n},$ (1)
where $\beta_{0}$ is the intercept, and each $\beta_{i}$ represents the effect of variable $X_{i}$ on the outcome. This structure is particularly useful as it allows researchers to include multiple predictors, control for potential confounders, and isolate the contribution of each variable to the outcome.
@@
Using a Bayesian approach provides a full posterior distributions over model parameters rather than point estimates, enabling direct uncertainty quantification and more robust inference, especially in settings with limited data, noisy measurements, or complex dependencies (for a foundational reference on Bayesian GLMs, see [McElreath, 2018]). This is particularly important in limited-data settings where confidence intervals based on the central limit theorem often underestimate uncertainty [Bowyer et al., 2025].
LLM evaluation data often exhibit structured dependencies such as multiple annotations from the same (or similar) graders, or items targeting related domains (e.g., several questions assessing historical knowledge). Our GLM framework naturally allows for hierarchical extensions that can account for these dependencies, enabling more robust inference and better-calibrated uncertainty [Luettgau et al., 2025].
Importantly, this framework supports answering the primary evaluation question (e.g., “how well does my LLM perform?”) while simultaneously identifying potential biases introduced by autograders. In the following sections we present illustrative examples of how these methods can be applied in practice.
@@
How are absolute scores modelled? Use GLMs for ordinal outcomes, such as ordered logistic regression Question 1
How are pairwise preferences modelled? Use GLMs for binary outcomes, such as binomial logistic regression Question 5
How do the scores of an autograder compare to those of a human expert? Include grader as a main effect Question 1.1
How can autograder reliability be assessed alongside model performance? Include both grader and LLM as predictors Question 1.2
How can self-bias be quantified? Include an interaction between grader and LLM; test for model-aligned preferences Question 2
How can human and autograder differences be analysed? Use a hierarchical GLM with grader-level effects nested in grader type (e.g., human vs. autograder) Question 3
Are some graders more lenient or strict than others? Estimate individual grader effects; inspect variation across graders Question 3
Do some items consistently receive higher or lower scores? Include item as a predictor; test whether some items are systematically easier or harder Question 4
Do graders disagree more on some items than others? Include grader x item interaction; test for grader-specific scoring patterns across questions Question 4
What is the uncertainty around inter-rater agreement metrics? Simulate scores from the model and compute agreement (e.g., Krippendorff’s $\alpha$) with uncertainty Question 4
Are longer responses favoured? Include token length (or token length difference) as a predictor Question 5
How can intransitive (e.g., cyclic) preferences be captured? Estimate pairwise probabilities and compare them across pairs Question 5
Is the grading scale well calibrated? Inspect cutpoints from ordered regression to analyse spacing and interpretability of score intervals appendix C
@@
Imagine a typical scenario: a researcher, Florence [^1] In tribute to the pioneering work of two Florence Nightingales in statistics: the 19th-century nurse who applied statistical methods to public health, and the 20th-century statistician who advanced combinatorics and statistical theory., is using LLMs to answer open-ended questions. She wants to assess how good her LLM is at this task, so decides to grade its answers. She develops a detailed rubric (table 2) with scores ranging from 1 (completely off-topic) to 10 (demonstrating deep understanding and insightful connections).
Florence carefully evaluates 100 examples using this rubric. For instance, when asked “Why was the initial World War II recommendation to reinforce areas of returning aircraft with bullet holes flawed?”, an answer stating only “Because they should have reinforced all areas equally rather than focusing on specific sections” might receive a score of 3, as it mentions a relevant idea (aircraft reinforcement strategy) but contains a fundamental misconception. The answer fails to recognise that the core issue was about misinterpreting the available data due to survivorship bias [Wald, 1939]. In contrast, a comprehensive response explaining that undamaged areas on returning aircraft actually represented the most vulnerable sections (as planes hit there did not survive to return), while also discussing the mathematical principles and broader implications of the survivor bias, might earn a score of 9 or 10.
@@
This manual evaluation process turns out to be very time-consuming, so Florence decides to switch to an autograder, providing it with her rubric and a few examples of graded responses. Outsourcing the grading of open-ended answers to an autograder is simply a means to make her assessment of the LLM’s question-answering capabilities faster and more scalable.
@@
This could have been the end of it, but being a careful researcher, she naturally asks herself, “Actually, can I trust my autograder?”. This crucial question leads to a lot of others, for example: “how will the scores of the autograder compare to my score?”, “Will the autograder favour generations from its own model family?”, “Will the autograder favour longer outputs?”. Using a GLM framework, she can formalise and answer these questions while simultaneously evaluating the LLM’s performance in answering open-ended questions.
@@
To do this, Florence needs to build a set of statistical models, select the one that best explains her data, and examine the estimated effects of different covariates (e.g., grader LLM, prompt, response length) and their interactions with the scores. For instance, by including a human (vs non-human) variable, she might discover that the autograder gives scores systematically 2 points higher than she does. Applying this approach across different covariates and interactions will allows her to understand what influences the scores, how to interpret results and what changes to make in the future.
@@
To illustrate how the framework can be applied in practice, we will follow Florence as she navigates common challenges in evaluating LLM outputs with autograders. All examples are based on simulated data that can be reproduced here. To encourage broader usage, all GLMs presented in this paper are implemented in the open-source HiBayes package and table 1 summarises the evaluation questions addressed, their corresponding GLM formulation and the sections where each is explained in detail.
@@ 1.1 Quantify the mean difference
As Florence wants to validate the autograder before using it, she decides to run it on the same N=100 examples that she had graded, and compare the results. In the left panel of fig. 1 we see a violin plot of the simulated scores given by Florence (human expert) in green, and by the autograder in yellow. Using our GLM framework we can fit a regression model with parameter estimates that quantifies the effect of the grader on the assigned score (eq. 2). Because the score can take values [1-10], we use an ordered logistic likelihood function, which accounts for the finite range and the categorical nature of the outcomes. This ordered logistic model has two main components: the linear predictor (cf. eq. 1) and the cutpoints. In our case, the linear predictor produces a continuous value on a latent (unobserved) scale. The ordered logistic model assumes that our discrete 1-10 scores are actually observations of this underlying continuous scale, divided by cutpoints. When the linear predictor falls between particular cutpoints $c_{j}$, we observe the corresponding discrete score. For example, if the linear predictor value falls between cutpoints $c_{3}$ and $c_{4}$, we would observe a score of 4. These cutpoints are estimated from the data along with the $\beta$ coefficients (see appendix C for more technical details).
@@
where $\phi_{i}$ is the linear predictor on the latent scale, $\boldsymbol{c}$ is the vector of cutpoints that divide the latent scale into discrete score categories, and the ordered logistic model uses a cumulative logit link function to map from the linear predictor to score probabilities through the relation: $\text{logit}(P(\text{score}_{i}\leq j))=c_{j}-\phi_{i}.$ The variable $X_{i}^{\text{grader}}$ encodes the identity of the grader who assigned score $i$ using effect coding [^2] Effect coding is a contrast coding scheme used in to represent categorical variables. Each level of the variable is assigned a numerical value such that the resulting coded variables are centered and sum to zero across all groups. This centering allows the intercept to reflect the grand mean (the average of the group means), and the coefficients represent deviations from that mean.. The linear model has two parameters: an intercept $\beta_{0}$, and a grader effect $\beta_{1}$. The intercept $\beta_{0}$ represents the average latent score across both graders, due to the centering imposed by effect coding. The grader coefficient $\beta_{1}$ captures the effect of the grader on the latent score and quantifies half the difference between the autograder and the human grader. Since the model encodes the autograder as $+1$ and the human as $-1$, the linear predictor shifts depending on the grader identity. These parameters define the location of the latent score, which is then mapped to the observed 1-10 categories via the threshold cutpoints. This shift in the latent score translates to a systematic difference in the probabilities of assigning higher or lower scores.
@@
To evaluate hypotheses, Florence fits different models to the data and compares their relative fit using statistical criteria such as Leave-One-Out cross-validation (LOO-CV) or the Widely Applicable Information Criterion (WAIC). For instance, here, Florence should compare a model that includes a grader effect (eq. 2) against a simpler model without this effect (i.e., $\phi_{i}=\beta_{0}$). The former statistical model tests the hypothesis that grader identity affects the assigned score, while the latter one serves as a null model with no grader effect. Model comparison indicates whether adding the grader effect improves model fit, but it does not guarantee that the grader parameter itself is significant. Therefore, once the best-fitting model (often called "winning model") is identified, Florence should inspect the parameter estimates and their credible intervals to confirm if there is clear statistical evidence for the grader effect.
@@
After fitting the statistical models and running a model comparison (fig. 7 in appendix B), we identify the model with the grader effect (eq. 2) as the best-fitting model. We can then examine the estimated effect of interest of this model (right panel of fig. 1). Since the grader variable is effect coded as $+1$ for the autograder and $-1$ for the human, the plotted difference in latent score between the autograder and the human grader is $2\beta_{1}$. This difference is defined as autograder minus human, so a negative value indicates that the human grader assigns higher scores on the latent scale. We observe that this difference is negative, with credible intervals that do not include zero, indicating a meaningful difference between the two graders. This confirms that the autograder systematically assigns lower scores than the human grader.

### v2

@@ Abstract
The evaluation of large language model (LLM) outputs is increasingly performed by other LLMs, a setup commonly known as “LLM-as-a-judge”, or autograders. While autograders offer a scalable alternative to human evaluation, they have shown mixed reliability and may exhibit systematic biases, depending on response type, scoring methodology, domain specificity, or other factors. Here we propose a statistical framework based on Bayesian generalised linear models (GLMs) that enables researchers to simultaneously assess their autograders while addressing their primary research questions (e.g., LLM evaluation). Our approach models evaluation outcomes (e.g., scores or pairwise preferences) as a function of properties of the grader (e.g., human vs. autograder) and the evaluated item (e.g., response length or the LLM that generated it), allowing for explicit quantification of scoring differences and potential biases within a unified framework. In addition, our method can be used to augment traditional metrics such as inter-rater agreement, by providing uncertainty estimates and clarifying sources of disagreement. Overall, this approach contributes to more robust and interpretable use of autograders in LLM evaluation, enabling both performance analysis and bias detection.
@@ Introduction
Large language models (LLMs) are being deployed across a wide range of tasks (e.g., code generation, question answering, decision support, autonomous tool use like calling APIs to retrieve documents). To understand their capabilities and ensure reliable performance and safety, LLM outputs need to be continuously and systematically evaluated. Given the high volume and increasing complexity of these evaluations, automated approaches have become common. Unsurprisingly, LLMs themselves have emerged as candidates for this task, a practice referred to as “LLM-as-a-judge” [Zheng et al., 2023; Liu et al., 2023; Chiang and Lee, 2023; Fu et al., 2024; Bavaresco et al., 2025; Wang et al., 2025] or as autograding [Schneider et al., 2023; Guan et al., 2024]. For brevity, we here refer to these models as autograders.
@@
Recent studies suggest that autograders may indeed exhibit systematic biases. One such bias is self-bias, where LLM-based graders tend to assign higher scores to responses generated by the same LLM [Panickssery et al., 2024; Liu et al., 2024b], or more broadly to machine-generated content over human-written responses [Liu et al., 2023]. Another common issue is length bias, where longer answers are preferred regardless of their actual quality [Zheng et al., 2023; Dubois et al., 2024]. Additional biases include preferences for certain writing styles, answer structures, or the presence of certain keywords [e.g., Koo et al., 2024; Wang et al., 2024; Stureborg et al., 2024; Wu and Aji, 2025]. Moreover, graders may exhibit intransitive preferences (e.g., preferring A over B, B over C, yet C over A), revealing inconsistencies that cannot be captured by standard grader evaluation metrics [Xu et al., 2025].
@@
Existing methods for autograder evaluation are limited in their ability to isolate and quantify such biases and inconsistencies. They do not explain why disagreement arises, do not account for how grader behaviour may vary with grader identity, model characteristics, or context, and do not quantify uncertainty in these effects. While recent work has started employing statistical modeling (e.g., logistic regression for length bias [Dubois et al., 2024]), these approaches typically remain narrowly tailored to specific biases and do not offer an integrated evaluation framework. In this paper, we introduce a Bayesian generalised linear modeling framework that complements existing metrics and provides an interpretable, extensible, and statistically grounded analysis of grader behaviour.
@@
In the following sections, we begin by presenting the general framework. We then walk through a series of common questions to demonstrate how the framework can be used to answer them. We begin by explaining how autograder scores can be compared to human scores (Question 1.1), and how this comparison can be integrated into an LLM evaluation analysis (Question 1.2). We then gradually increase the complexity of the setup and show how different questions about autograder quality can be addressed within this integrated framework. Specifically, we show how to assess whether the graders are biased towards certain models being evaluated (Question 2) and how to quantify individual differences in the case of multiple graders (Question 3). We also show how to analyse item-level patterns (Question 4), which allows us to compute inter-rater agreement metrics with quantification of uncertainty and, crucially, to identify whether disagreement stems from noise or systematic bias (something that traditional metrics cannot disentangle). Finally, we show how this framework can be applied to pairwise judgments settings, in which LLM outputs are compared directly to one another (as opposed to being assigned absolute scores), how to quantify intransitive (e.g., cyclic) preferences and how to assess whether graders have biases towards longer formats (Question 5). In table 1, we provide a summary of the evaluation questions along with their corresponding formalisations as a practical guide for researchers using our framework.
@@
We propose using generalised linear models (GLMs) as a practical and flexible approach for analysing autograder performance during LLM evaluations. GLMs extend standard linear regression by allowing the outcome variable to follow a wider range of distributions, making them suitable for many data types beyond normally distributed outcomes. At their core, GLMs relate the expected outcome to a linear predictor via a link function:
$g(\mu)=\beta_{0}+\beta_{1}X_{1}+\beta_{2}X_{2}+\dots+\beta_{n}X_{n}$ (1)
where $\mu=\text{E}[Y]$ represents the expected outcome, $g(\cdot)$ is a link function appropriate to the distribution of the outcome, $\beta_{0}$ is the intercept, and each $\beta_{i}$ represents the effect of predictor variable $X_{i}$. This structure is particularly useful as it allows researchers to include multiple predictors, control for potential confounders, and isolate the contribution of each variable to the outcome.
@@
Using a Bayesian approach provides full posterior distributions over model parameters rather than point estimates, enabling direct uncertainty quantification and more robust inference, especially in settings with limited data, noisy measurements, or complex dependencies (for a foundational reference on Bayesian GLMs, see McElreath [2018]). This is particularly important in limited-data settings where confidence intervals based on the central limit theorem often underestimate uncertainty [Bowyer et al., 2025].
Moreover, LLM evaluation data often exhibit structured dependencies such as multiple annotations from the same (or similar) graders, or items targeting related domains (e.g., several questions assessing historical knowledge). GLMs naturally allow for hierarchical extensions that can account for these dependencies, enabling more robust inference and better-calibrated uncertainty [Luettgau et al., 2025].
Importantly, a GLM-based framework supports answering the primary evaluation question (e.g., “how well does my LLM perform?”) while simultaneously identifying potential biases introduced by autograders. In the following sections we present illustrative examples of how these methods can be applied in practice.
@@
What grades does my autograder give compared to a human grader? Include grader as a main effect. Question 1.1
Can I assess the quality of my autograder during my evaluations? Include both grader and LLM as predictors. Question 1.2
Do my autograder(s) favour their own output? Include an interaction between grader and LLM; test for model-aligned preferences. Question 2
Is there a general human vs autograder difference? Use a hierarchical GLM with grader-level effects nested in grader type (e.g., human vs. autograder). Question 3
Are some graders more lenient or strict than others? Estimate individual grader effects; inspect variation across graders. Question 3
Do some items consistently receive higher or lower scores? Include item as a predictor; test whether some items are systematically easier or harder. Question 4
Do graders disagree more on some items than others? Include grader x item interaction; test for grader-specific scoring patterns across questions. Question 4
What is the uncertainty around inter-rater agreement metrics? Simulate scores from the model and compute agreement (e.g., Krippendorff’s $\alpha$) with uncertainty. Question 4
Do grader(s) favour longer responses? Include token length (or token length difference) as a predictor. Question 5
Do my grader(s) exhibit intransitive (e.g., cyclic) preferences? Estimate pairwise probabilities and compare them across pairs. Question 5
Is my grading scale well calibrated? Inspect cutpoints from ordered regression to analyse spacing and interpretability of score intervals. appendix C
@@
Imagine a typical scenario: a researcher, Florence [^1] In tribute to the pioneering work of two Florence Nightingales in statistics: the 19th-century nurse who applied statistical methods to public health and the 20th-century statistician who advanced combinatorics and statistical theory., is using LLMs to answer open-ended questions. To evaluate the quality of the model’s responses, she decides to manually grade them. She develops a detailed rubric (table 2) with scores ranging from 1 (completely off-topic) to 10 (demonstrating deep understanding and insightful connections).
Florence carefully evaluates N=100 model responses using her rubric. For example, consider the question: “During World War II, which part of aircraft did statisticians decide to reinforce?”. A low-scoring response (e.g., 3 points) might simply say: “The areas with the least bullet holes.” While this demonstrates some understanding, it misses the important statistical reasoning behind this strategy. In contrast, a high-scoring response (e.g., 9 or 10 points) would explain that they focused on areas with no damage, because aircraft hit in those places often did not survive to return. The response might further elaborate on the statistical principle of survivorship bias [Wald, 1939] and its practical implications.
@@
This manual evaluation process proves to be very time-consuming, so Florence decides to turn to an autograder. She provides it with her rubric and a few examples of graded responses. By outsourcing the grading of open-ended answers to an autograder, she makes the assessment of the LLM’s question-answering capabilities not only more efficient but also significantly more scalable.
@@
This could have been the end of it, but being a careful researcher, she naturally asks herself, “Can I actually trust my autograder?”. This central question unfolds into several more specific ones, for example: “How do the autograder’s scores compare to mine?”, “Does the autograder favour generations produced by LLMs from its own model family?”, “Does it tend to reward longer answers?”. Using a GLM framework, she can formalise and answer these questions while simultaneously evaluating the LLM’s performance in answering open-ended questions.
@@
To do this, Florence needs to build a set of statistical models, select the one that best explains her data, and examine how different covariates (e.g., grader, LLM, prompt, response length) and their interactions among one another systematically influence the scores. For instance, by including a human (vs. non-human) variable, she might discover that the autograder gives scores systematically 2 points higher than she does. Applying this approach across different covariates and interactions allows her to identify specific biases (such as model favouritism or length preferences), understand how grading patterns arise, and make informed adjustments to the autograder or evaluation process.
@@
To illustrate how the framework can be applied in practice, we will follow Florence as she navigates common challenges in evaluating LLM outputs with autograders. All examples are based on simulated data, which can be reproduced from a publicly available repository. To facilitate wider adoption, all GLMs presented in this paper are implemented in the open-source HiBayes package. table 1 summarises the evaluation questions addressed, their corresponding GLM formulation, and the sections where each is presented and explained in detail.
@@ 1.1 Quantify the mean difference
As Florence wants to validate the autograder before using it, she decides to run it on the same N=100 examples that she had graded, and compare the results. In the left panel of fig. 1, we see a violin plot of the simulated scores given by Florence (human expert) in green, and by the autograder in yellow. Using our GLM framework we can fit a regression model with parameter estimates that quantifies the effect of the grader on the assigned score (eq. 2). Because the score can take values [1-10], we use an ordered logistic likelihood function, which accounts for the finite range and the categorical nature of the outcomes. This ordered logistic model has two main components: the linear predictor (cf. eq. 1) and the cutpoints. In our case, the linear predictor produces a continuous value on a latent (unobserved) scale. The ordered logistic model assumes that our discrete 1-10 scores are actually observations of this underlying continuous scale, divided by cutpoints. When the linear predictor falls between particular cutpoints $c_{j}$, we observe the corresponding discrete score. For example, if the linear predictor value falls between cutpoints $c_{3}$ and $c_{4}$, we would observe a score of 4. These cutpoints are estimated from the data along with the $\beta$ coefficients (see appendix C for more technical details).
@@
where $\phi_{i}$ is the linear predictor on the latent scale, $\boldsymbol{c}$ is the vector of cutpoints that divide the latent scale into discrete score categories, and the ordered logistic model uses a cumulative logit link function to map from the linear predictor to score probabilities through the relation: $\text{logit}(P(\text{score}_{i}\leq j))=c_{j}-\phi_{i}.$ The variable $X_{i}^{\text{grader}}$ encodes the identity of the grader who assigned score $i$ using effect coding. [^2] Effect coding is a contrast coding scheme used in to represent categorical variables. Each level of the variable is assigned a numerical value such that the resulting coded variables are centered and sum to zero across all groups. This centering allows the intercept to reflect the grand mean (the average of the group means), and the coefficients represent deviations from that mean. The linear model has two parameters: an intercept $\beta_{0}$, and a grader effect $\beta_{1}$. The intercept $\beta_{0}$ represents the average latent score across both graders, due to the centering imposed by effect coding. The grader coefficient $\beta_{1}$ captures the effect of the grader on the latent score and quantifies half the difference between the autograder and the human grader. Since the model encodes the autograder as $+1$ and the human as $-1$, the linear predictor shifts depending on the grader identity. These parameters define the location of the latent score, which is then mapped to the observed 1-10 categories via the threshold cutpoints. This shift in the latent score translates to a systematic difference in the probabilities of assigning higher or lower scores.
@@
To evaluate hypotheses, Florence fits different models to the data and compares their relative fit using statistical criteria such as Leave-One-Out cross-validation (LOO-CV) or the Widely Applicable Information Criterion (WAIC). For instance, here, Florence should compare a model that includes a grader effect (eq. 2) against a simpler model without this effect (i.e., $\phi_{i}=\beta_{0}$). The former statistical model tests the hypothesis that grader identity affects the assigned score, while the latter one serves as a null model with no grader effect. Model comparison indicates whether adding the grader effect improves model fit, but it does not guarantee that the grader parameter itself is significant. Therefore, once the best-fitting model (often called “winning model”) is identified, Florence should inspect the parameter estimates and their credible intervals to confirm if there is clear statistical evidence for the grader effect.
@@
After fitting the statistical models and running a model comparison (fig. 8 in appendix B), we identify the model with the grader effect (eq. 2) as the best-fitting model. We can then examine the estimated effect of interest of this model (right panel of fig. 1). Since the grader variable is effect coded as $+1$ for the autograder and $-1$ for the human, the plotted difference in latent score between the autograder and the human grader is $2\beta_{1}$. This difference is defined as autograder minus human, so a negative value indicates that the human grader assigns higher scores on the latent scale. We observe a negative difference, with credible intervals that do not include zero, indicating a meaningful difference between the two graders. This confirms that the autograder systematically assigns lower scores than the human grader.